# The evolution of inharmonicity and noisiness in contemporary popular music

### v1

@@ 6.1 Conventions
    The term partial denotes the spectral representation of a single periodic sine wave.
@@
    A harmonic complex tone designates an ensemble of partials that are integer multiples of a fundamental frequency.
@@
    An inharmonic complex tone designates a tone (i.e., it is perceived to have a pitch) that consists of an ensemble of partials that are near-integer multiples of a fundamental frequency. Fletcher et al., (1962) and Rasch and Plomp, (1982, p. 9) provide examples of such tones.
@@
    In a harmonic or inharmonic complex tone, an overtone can be any of the partials except the one that corresponds to the fundamental frequency.
@@
    A harmonic is any sinusoidal component of a harmonic complex tone, with the $n$th harmonic having a frequency that is $n$ times the frequency of the fundamental (so the fundamental is the first harmonic).
In this section, harmonic complex tones are generated using the model mentioned in Mauch and Dixon, (2010) and previously introduced by Gómez, (2006), according to which the $k$th partial, denoted by $a(k)$, is assigned an amplitude $s^{k-1}$ for some constant $s<1$. Unless stated otherwise, we use 10 harmonics over a 220Hz fundamental (A3) and set $s=0.8$.
6.2 One complex tone
We start with the case of a single isolated complex tone. Figure 9 shows the evolution of HarmonicRatio values in four cases as follows.
@@ 6.2 One complex tone
    Increasing number of inharmonic partials in an initially harmonic tone.
@@
    Increasing number of inharmonic partials in an initially harmonic tone, with harmonics having less energy than in case 1 ($s=0.7$).
@@
    Frequency shifting of one single partial (arbitrarily, partial 4) of an initially harmonic tone.
@@
    Superposition of two sine waves of equal power. The frequency of the first sine wave is 220Hz. The frequency of the second sine wave ranges from 220Hz to 880Hz.
Figure 9: (a) HarmonicRatio resulting from an increasing number of inharmonic partials in an initially harmonic complex tone ($f0=200$Hz, 10 harmonics). The inharmonicity of the partials is random, the representation summarises observations from 50 tones, the black line denotes the median over the 50 observations, and the gray area the 25th and 75th percentiles. (b) Same as top graph, but with harmonics of lower energy. (c) HarmonicRatio resulting from the frequency shift of a single partial. (d) HarmonicRatio resulting from the superposition of two sine waves depending on the interval between them. The text boxes show the frequency ratios of the ‘purer’ intervals inside the first octave. The $y$-axis scale is identical for all graphs in order to facilitate comparison.
The results are shown in the four graphs in Figure 9, from which we can derive the following observations:
@@
    Figure 9 (a) confirms the MPEG-7 specification’s interpretation (ISO,, 2001), in the sense that an increasing proportion of inharmonic components in an initially harmonic complex tone indeed decreases the HarmonicRatio value.
@@
    Figure 9 (b) shows that the resulting HarmonicRatio value depends on the energy of the inharmonic components. As the inharmonic components become relatively more intense, the HarmonicRatio value falls.
@@
    Figure 9 (c) shows that when one single partial is inharmonic, the resulting HarmonicRatio value depends on the partial’s frequency.
@@
    Figure 9 (d) elaborates on the influence of the intervals between two partials on the resulting HarmonicRatio values. (i) ‘Purer’ or ‘simpler’ intervals, i.e., intervals that can be expressed as a fraction composed of smaller integers (Vos and van Vianen,, 1985), are local maxima of the HarmonicRatio values. The ratio’s denominator is a key influence on the HarmonicRatio of the interval. (ii) Intervals result in comparatively higher HarmonicRatio values as they get wider. Observations (i) and (ii) are reminiscent of sensory dissonance (Sethares,, 2005, p.46, p.85), and what Masina et al., (2022) refer to as ‘roughness’, following Hutchinson and Knopoff, (1978) and Plomp and Levelt, (1965).
@@
6.3 Several complex tones
The properties leading to low HarmonicRatio values in single complex tones remain valid in the case of the superposition of several complex tones. One difference lies in the fact that the distribution of the complex tones’ fundamentals often reflects scales or tuning systems (e.g., a diatonic scale or 12-tone equal temperament). We examine the HarmonicRatio values resulting from the superposition of complex tones whose fundamentals conform to different scales. We start from one harmonic complex tone, and add complex tones whose fundamentals range from 220Hz to 2200Hz. We compare the HarmonicRatio measures in four cases:
@@ 6.3 Several complex tones
    A higher number of harmonic tones leads to decreasing HarmonicRatio values. In other words, the more voices in the music, the more it becomes inharmonic.
@@
    HarmonicRatio values are influenced by the scales on which the music relies. HarmonicRatio is lowest when the scale is continuous (case 1). It increases when the scale is restricted to the twelve-tone chromatic scale (case 2). It increases again when the fundamentals only constitute major triads (case 3). Finally, it remains similar when the major triads use pure intervals instead of equal-tempered intervals (case 4).
@@
    The very low HarmonicRatio values resulting from fundamental frequencies not following a discrete scale suggest that the presence of continuous frequency contours (e.g., human voices in music) may be a factor that lowers HarmonicRatio values.
Additionally, two factors that can influence HarmonicRatio values in the case of several complex tones derive from observations that were made in section 6.2 in the case of one single complex tone:
@@
    Proximity of f0s. We saw that closer partials generally result in lower HarmonicRatio values. As a result, fundamental frequencies that are close to each other (as may happen, for instance, in near-coincident musical voices) will result in relatively low HarmonicRatio values.
@@
    Voices of equal energies. In the case where all partials are inharmonic, e.g. in piano sounds (Fletcher et al.,, 1962), the lowest HarmonicRatio values will be obtained when the partials have approximately equal energy. This may happen with two musical voices of close energy. If the fundamentals are close in frequency, then the lowest HarmonicRatio values will be obtained if the fundamentals have the same energy.
To summarise, given several complex tones, low HarmonicRatio values may be the combined result of
@@
    a continuous distribution of f0 values.
6.4 Respective contribution of fundamental and overtones
Figure 11: First row, HarmonicRatio derived from the pairwise combination of elements in two harmonic tones. Clearer shades correspond to higher values, and darker shades to lower values. The interval between the two tones is (a) 3 half-tones and (b) 5 half-tones. Partial numbers are written as ‘tone number-partial number’ (e.g., ‘2-3’ corresponds to the third partial of the second tone). Second row, (c) and (d), HarmonicRatio derived from the pairwise combination of elements in two inharmonic tones (same fundamentals as first row, inharmonicity coefficient is 0.002 as defined by Fletcher et al., (1962)).
Figure 11, graphs (a) and (b) show the HarmonicRatio values resulting from the pairwise superposition of each component of two harmonic complex tones. The tones leading to Figure 11 (a) and (c) have fundamentals that are 3 semitones apart (an equal-tempered minor third). The tones leading to Figure 11 (b) and (d) have fundamentals that are 5 semitones apart (an equal-tempered perfect fourth).
As shown in Figure 9 (d) in the case of two sine waves, the equal-tempered minor third corresponds to a lower HarmonicRatio than the perfect fourth. In the case of complex tones, we can make a distinction between harmonic and inharmonic complex tones:
@@ 6.4 Respective contribution of fundamental and overtones
    Two harmonic complex tones (the two top graphs in Fig. 11)    For both the minor third and the perfect fourth, HarmonicRatio values deriving from the superposition of overtones inside one harmonic complex tone are higher than HarmonicRatio deriving from the superposition of overtones belonging to different harmonic complex tones. The observation would be valid for all intervals except the octave, in which case all the harmonics of the upper tone would coincide with harmonics of the lower tone and the HarmonicRatio values resulting from the interaction between harmonics within one of the complex tones would be no different from those resulting from the interaction between harmonics in different tones.
@@
    Two inharmonic complex tones (the two bottom graphs in Fig. 11)    HarmonicRatio values resulting from the superposition of partials within a complex tone are generally lower than in the case of harmonic complex tones. In the case of the five-semitone interval (right-hand side of figure), they are lower than the values resulting from the superposition of the fundamentals. The observation would be valid for all intervals corresponding to a high HarmonicRatio value.
@@
    Increasing the inharmonicity of individual complex tones reduces the relative influence on inharmonicity of the intervals between complex tones.
6.5 HarmonicRatio and acoustic beating
As mentioned in section 6.2, the behaviour shown by the graph at the bottom of Figure 9 is reminiscent of that of ‘roughness’. According to Masina et al., (2022), relations between the sensation of ‘roughness’ and acoustic beats were introduced in 1832–1837 by Foderà (Barbieri,, 2002) and in the middle of the 19th century by Helmholtz (Helmholtz and Ellis,, 1885, pp. 197-211). Rasch and Plomp, (1982, p. 15) illustrate how acoustic beating corresponds to the sensation of ‘roughness’ when the interval between two tones is smaller than that of a critical band. We compare measures of acoustic beating and HarmonicRatio values deriving from the superposition of harmonic complex tones.
@@ 6.5 HarmonicRatio and acoustic beating
Figure 12 (a) is the result of the same process as Figure 9 (d) using two harmonic complex tones instead of two sine waves. The f0 of the first complex tone is 220Hz and the f0 of the second complex tone ranges from 220Hz to 880Hz. In Figure 12 (b), we superpose the HarmonicRatio values from Figure 12, top, with measures of acoustic beating (background image). We evaluate acoustic beating by computing the power spectrum of the envelope (0.05s windows) of the sum of the two complex tones. High power spectrum values correspond to the presence of acoustic beating at the corresponding frequency. We use a higher number of harmonics (40) and a higher $s$ factor (0.95) so that the white lines deriving from acoustic beating are more visible. From these results, we can make the following observations:
@@
    Pure intervals correspond to infinitely slow acoustic beating (i.e., no beating). Increased deviation from pure intervals corresponds to acoustic beating becoming faster. In that sense, more ‘roughness’ may correspond to faster acoustic beats.
@@
    The local maxima of the HarmonicRatio values occur when acoustic beating is slowest. In that sense, high HarmonicRatio values is associated with low ‘roughness’ and slow acoustic beating.
From this point of view, results involving HarmonicRatio (such as in section 4) may equally apply to the period of the acoustic beating, in other words, the ‘roughness’ or acoustical ‘grit’ of the music.
6.6 HarmonicRatio and noise
The tones used to produce the results shown in Figures 912 were each composed of a finite number of sinusoidal components. Such signals do not contain noise. As previously mentioned in reference to Schneider and Frieler, (2009), inharmonicity can also stem from noise. In this case, low HarmonicRatio values can be understood in principle as being the result of the presence of inharmonic relations between the (infinite set of notional) sine waves that constitute the noise whose frequencies can be infinitesimally different from each other.
From the perspective of discrete spectral transforms, an infinite-length signal can be identified as devoid of noise when spectral bins with non-zero values are separated by bins with zero values. Conversely, the presence of noise may be characterised by several contiguous bins with non-zero values. In practice, however, it is difficult to define precisely what constitutes ‘noise’ in a spectral transform because
@@ 6.6 HarmonicRatio and noise
    even the spectra of harmonic tones are not actually line spectra because they do not have infinite duration; and
@@
    a transform takes place on an audio window that covers a time span so that a sine wave with even small frequency changes within the window (such as the result of a vibrato) will not result in a line but in a distribution of frequencies.
@@
As it is difficult to distinguish noise from spectral lines in practice, we choose not to distinguish between them. We fall back on the objective information we have. Thus, from a spectral domain perspective, the signal’s information is represented as values attached to spectral bins. Frequency intervals are only defined between the spectral bin center frequencies, and any information we derive from the signal originates from the values attached to the spectral bins. From this point of view, a feature representing the ‘degree of inharmonicity’ in the signal may be conceptualised as deriving from the ensemble of pairwise combinations between all frequency bins (positions and energy values). As a result, a feature representing the ‘degree of inharmonicity’ in the signal will not be able to distinguish between ‘inharmonic sounds which have little if any relevance for music (e.g., white or pink noise)’ (Schneider and Frieler,, 2009) and ‘coherent’ inharmonic signals, which ‘sound as stable and smooth as harmonic signals’ (de Boer,, 1956), i.e., inharmonic complex tones with a finite number of sinusoidal components. In light of the relation between HarmonicRatio and pitch strength mentioned at the beginning of section 3.2.1, the ‘first peak of the auto-correlation function’ (Patterson et al.,, 1996; Yost,, 1996, 1997; Shofner and Selas,, 2002) is unable to determine whether a lower pitch strength perception derives from relations between partials, or from noise. We need to find a method to estimate the relative extent to which inharmonicity results from the interaction between (approximately) discrete partials as opposed to the presence of noise.
@@
Figure 13: HarmonicRatio, inharmonicity and noisy signals. (a) The signals corresponding to the two spectra have the same HarmonicRatio value of 0.9558. The red signal is inharmonic, with an inharmonicity coefficient of 0.002 according to Fletcher et al., (1962). The blue signal is the sum of a harmonic complex tone and white noise. The white noise’s RMS power is 13.5dB below the harmonic tone’s. (b) Comparative decrease of HarmonicRatio values starting from a sine wave. The blue curve is the result of added white noise. The red curve is the result of adding sine waves of random frequencies. The frequencies being random, the experiment is repeated 200 times. The red area shows the 25th and 75th percentiles of the results.
@@
Figure 13 (a) illustrates how an inharmonic complex tone and a harmonic complex tone with white noise can result in the same HarmonicRatio value. Conversely, given a HarmonicRatio value less than 1, the original audio may consist of a non-inharmonic tone superposed with noise, or of a noiseless inharmonic complex tone.
@@
Figure 13 (b) illustrates how a sum of sine waves and a sine wave with white noise can result in the same HarmonicRatio values. Specifically, in this graph it can be seen that a sine wave superposed with noise of equal RMS power is as inharmonic on average as a sum of 14 sine waves of the same energy and random frequencies.
6.7 Inharmonicity as 1--HarmonicRatio.
@@
The output of the HarmonicRatio feature can be viewed as a reflection of the pairwise relations between the values in all spectral bins, regardless of whether the spectral bin values originate from noise or from partials. High HarmonicRatio values correspond to harmonic partials without noise. Any deviation will lower the resulting HarmonicRatio values. The deviation can derive from inharmonicity in complex tones, inharmonic relations that derive from certain combinations of complex tones, and inharmonic relations deriving from the presence of contiguous non-zero spectral bins, which can originate from noise. This summary characterises an ‘inharmonicity’ feature, which we call ‘HR-inharmonicity’ and directly derive from HarmonicRatio:
@@
7 ‘Noisiness’, spectral flatness and peak prominence
In the previous section, we established that:
@@
    A harmonic complex tone with noise and a noiseless inharmonic complex tone may lead to the same value of HarmonicRatio.
@@
    A discrete ensemble of sine waves with no harmonic relations and white noise added to a sine wave may lead to the same value of HarmonicRatio.
@@
In order to discriminate between noiseless and noisy signals leading to the same value of HarmonicRatio, we need a feature that measures the noisiness of the signal. To that end, we introduce a new metric derived from the spectral flatness (Peeters,, 2004) and AudioSpectrumFlatnessType (ISO,, 2001) features, which is robust to the overall spectral profile.
@@
Peeters, (2004, p. 20) defines spectral flatness as ‘a measure of the noisiness […] / sinusoidality of a spectrum’. It is computed using the ratio of the geometric mean to the arithmetic mean, which is equivalent to the Wiener entropy (WE) of the energy spectrum value. A comparison between the WE and the variance of the power spectra derived from the BEA dataset shows that the two measures are highly correlated (Pearson correlation = 0.9973). As a result, in the context of this study, it is possible to intuitively understand spectral flatness as similar to the variance of the power spectrum. Intuitively, low spectral flatness values (i.e., high WE and high variance) denote the presence of spectral peaks, while high values (i.e., low WE and low variance, ‘flat’ spectrum) denote a noise-like signal.
@@
Spectral flatness is included in the Essentia toolbox (Bogdanov et al.,, 2013) as FlatnessDB [^11] https://essentia.upf.edu/reference/std_FlatnessDB.html and in the MPEG-7 standard as AudioSpectrumFlatnessType (ISO,, 2001, p. 26). For the evaluation of the spectrum, the MPEG-7 documentation specifies a logarithmic frequency resolution of a 1/4 octave. We prefer to use a higher resolution so that the spectral peaks are more clearly defined. We use one-quarter of a semi-tone (25 cents). This interval is close to a Pythagorean comma and of roughly the same order of magnitude as the minimum perceivable pitch difference (Zarate et al.,, 2012) or ‘just noticeable difference’ (Stern and Johnson,, 2010) for pitch. The use of such a higher resolution is made possible by the analysis taking place on the Fourier transform instead of the short-term Fourier transform as is the case in the MPEG-7 standard.
@@
As a preliminary step, we calculate spectral flatness for the music datasets, for pink noise, and for white noise. The output value is $0.998$ for white noise. The value for pink noise ($0.84$) is between the values for the BEA dataset ($0.83$) and the orchestra dataset ($0.86$). The difference between the values for white noise and pink noise stems from the fact that the original spectrum is evaluated on a logarithmic frequency scale. In the case of white noise, bands are of constant width in terms of frequency; whereas, in the case of pink noise, they are decreasing. When we use a linear frequency scale, the situation is reversed. Spectral flatness is sensitive to the global envelope of the spectrum. The fact that the spectral flatness values of pink noise and music are similar, defeats the original purpose of measuring noisiness. We want to measure the prominence of the peaks, in other words, the noisiness of the signal, independently of the global envelope. Therefore, we perform median filtering on the spectrum before evaluating its WE. The filtering involves 8-bin windows (1 tone). Experiments show that window widths between 4 and 16 bins lead to comparable results.

### v2

@@ A.1.1 Conventions
    The term partial refers to the spectral representation of a single periodic sine wave.
@@
    A harmonic complex tone denotes a group of partials that are integer multiples of a fundamental frequency.
@@
    An inharmonic complex tone refers to a tone (perceived to have a pitch) consisting of partials that are near-integer multiples of a fundamental frequency. Examples are given by Fletcher et al., (1962) and Rasch and Plomp, (1982, p. 9).
@@
    In harmonic or inharmonic complex tones, an overtone is any partial except the one corresponding to the fundamental frequency.
@@
    A harmonic is any sinusoidal component of a harmonic complex tone, with the $n$th harmonic having a frequency $n$ times that of the fundamental (the first harmonic).
In this section, harmonic complex tones are generated using the model from Mauch and Dixon, (2010), originally introduced by Gómez, (2006), where the $k$th partial, $a(k)$, is assigned an amplitude of $s^{k-1}$ for a constant $s<1$. Unless specified otherwise, we use 10 harmonics with a 220Hz fundamental (A3) and set $s=0.8$.
A.1.2 One complex tone
We begin with the case of a single isolated complex tone. Figure 16 shows the evolution of HarmonicRatio values in four cases:
@@ A.1.2 One complex tone
    Increasing the number of inharmonic partials in an initially harmonic tone.
@@
    Increasing the number of inharmonic partials in an initially harmonic tone, with harmonics having less energy than in case 1 ($s=0.7$).
@@
    Frequency shift of a single partial (here, partial 4) in an initially harmonic tone.
@@
    Superposition of two sine waves of equal power. The first sine wave has a frequency of 220Hz, while the second ranges from 220Hz to 880Hz.
Figure 16: (a) HarmonicRatio resulting from an increasing number of inharmonic partials in an initially harmonic complex tone ($f_{0}=200$Hz, 10 harmonics). The inharmonicity of the partials is random, the representation summarises observations from 50 tones, the black line denotes the median over the 50 observations, and the gray area the 25th and 75th percentiles. (b) Same as top graph, but with harmonics of lower energy. (c) HarmonicRatio resulting from the frequency shift of a single partial. (d) HarmonicRatio resulting from the superposition of two sine waves depending on the interval between them. The text boxes show the frequency ratios of the ‘purer’ intervals inside the first octave. The $y$-axis scale is identical for all graphs in order to facilitate comparison.
The results are shown in Figure 16, from which we can derive the following observations:
@@
    Figure 16 (a) confirms the MPEG-7 specification’s interpretation (ISO,, 2001), showing that increasing inharmonic components in an initially harmonic complex tone decreases the HarmonicRatio value.
@@
    Figure 16 (b) shows that HarmonicRatio depends on the energy of the inharmonic components. As these components grow more intense, the HarmonicRatio value drops.
@@
    Figure 16 (c) shows that when one partial is inharmonic, HarmonicRatio depends on the partial’s frequency.
@@
    Figure 16 (d) shows the influence of intervals between two partials on HarmonicRatio. (i) Simpler intervals, i.e., fractions with smaller integers (Vos and van Vianen,, 1985), are local maxima. The denominator influences the interval’s HarmonicRatio. (ii) Wider intervals yield higher HarmonicRatio values. These findings recall sensory dissonance (Sethares,, 2005, p.46, p.85), and what Masina et al., (2022) refer to as ‘roughness’, following Hutchinson and Knopoff, (1978) and Plomp and Levelt, (1965).
@@
A.1.3 Several complex tones
The factors leading to low HarmonicRatio values in single complex tones also apply when several complex tones are superposed. One difference is that the distribution of the complex tones’ fundamentals often reflects scales or tuning systems (e.g., diatonic scale or 12-tone equal temperament). We explore the HarmonicRatio values arising from the superposition of complex tones with fundamentals conforming to different scales. Starting with one harmonic complex tone, we add others with fundamentals ranging from 220Hz to 2200Hz. We compare the HarmonicRatio measures in four cases:
@@ A.1.3 Several complex tones
    A higher number of harmonic tones leads to lower HarmonicRatio values. In other words, more voices in the music increase inharmonicity.
@@
    HarmonicRatio values are affected by the scales used in the music. HarmonicRatio is lowest for a continuous scale (case 1). It increases for the twelve-tone chromatic scale (case 2), rises further with major triads (case 3), and remains similar when pure intervals replace equal-tempered intervals in major triads (case 4).
@@
    Very low HarmonicRatio values from fundamentals not following a discrete scale suggest that continuous frequency contours (e.g., human voices) may lower HarmonicRatio values.
Two factors influencing HarmonicRatio values for multiple complex tones stem from the observations in Appendix A.1.2 for a single complex tone:
@@
    Proximity of $f_{0}$s. Closer partials tend to lower HarmonicRatio values. Thus, nearby fundamental frequencies (e.g., in near-coincident musical voices) lead to lower HarmonicRatio values.
@@
    Voices of equal energies. When all partials are inharmonic (e.g., piano sounds (Fletcher et al.,, 1962)), the lowest HarmonicRatio values occur when partials have equal energy. If the fundamentals are close in frequency, the lowest HarmonicRatio values arise when they share equal energy.
To summarise, given several complex tones, low HarmonicRatio values may be the combined result of:
@@
    a continuous distribution of $f_{0}$ values.
A.1.4 Respective contribution of fundamental and overtones
Figure 18 (a) and (b) show the HarmonicRatio values from the pairwise superposition of components of two harmonic complex tones. The tones in Figure 18 (a) and (c) have fundamentals 3 semitones apart (an equal-tempered minor third), while those in Figure 18 (b) and (d) have fundamentals 5 semitones apart (an equal-tempered perfect fourth).
As shown in Figure 16 (d), for two sine waves, the equal-tempered minor third yields a lower HarmonicRatio than the perfect fourth. In the case of complex tones, we can distinguish between harmonic and inharmonic complex tones:
@@ A.1.4 Respective contribution of fundamental and overtones
    Two harmonic complex tones (Figure 18 (a) and (b))    For both the minor third and the perfect fourth, HarmonicRatio values from the superposition of overtones within a single harmonic complex tone are higher than those from the superposition of overtones between different harmonic complex tones. This applies to all intervals except the octave, where all harmonics of the upper tone coincide with those of the lower tone, making the HarmonicRatio values from interactions within one tone the same as those between different tones.
    Figure 18: (a) and (b) show HarmonicRatio values derived from the pairwise combination of elements in two harmonic tones. Clearer shades correspond to higher values, and darker shades to lower values. The interval between the two tones is 3 half-tones in (a) and 5 half-tones in (b). Partial numbers are written as ‘tone number-partial number’ (e.g., ‘2-3’ corresponds to the third partial of the second tone). (c) and (d) show HarmonicRatio values derived from the pairwise combination of elements in two inharmonic tones (same fundamentals as in (a) and (b), respectively). The inharmonicity coefficient in (c) and (d) is 0.002 as defined by Fletcher et al., (1962).
@@
    Two inharmonic complex tones (Figure 18 (c) and (d))    HarmonicRatio values from the superposition of partials within the tones are generally lower than for harmonic complex tones. For the five-semitone interval, they are lower than those from the superposition of the fundamentals. This observation would hold for all intervals with a high HarmonicRatio value.
@@
    Increasing the inharmonicity of individual complex tones reduces the relative influence on the inharmonicity of the intervals between complex tones.
A.1.5 HarmonicRatio and acoustic beating
As noted in Appendix A.1.2, the behaviour in the bottom graph of Figure 16 resembles that of ‘roughness’. According to Masina et al., (2022), the relation between ‘roughness’ and acoustic beats was introduced by Foderà in 1832–1837 (Barbieri,, 2002) and later developed by Helmholtz in the 19th century (Helmholtz,, 1885, pp. 197-211). Rasch and Plomp, (1982, p. 15) explain how acoustic beating corresponds to ‘roughness’ when the interval between two tones is smaller than a critical band. We compare measures of acoustic beating with HarmonicRatio values from the superposition of harmonic complex tones.
@@ A.1.5 HarmonicRatio and acoustic beating
Figure 19 (a) follows the same process as Figure 16 (d) but uses two harmonic complex tones instead of two sine waves. The $f_{0}$ of the first tone is 220Hz, while the second ranges from 220Hz to 880Hz. In Figure 19 (b), we superpose the HarmonicRatio values from Figure 19 (a) with acoustic beating measures (background image). Acoustic beating is evaluated by computing the power spectrum of the envelope (0.05s windows) of the sum of the two tones. High power spectrum values indicate the presence of acoustic beating at that frequency. A higher number of harmonics (40) and a higher $s$ factor (0.95) are used to enhance the visibility of the white lines from acoustic beating. From these results, we make the following observations:
@@
    Pure intervals correspond to infinitely slow acoustic beating (i.e., no beating). Greater deviation from pure intervals leads to faster acoustic beating. Thus, increased ‘roughness’ may correspond to faster acoustic beats.
@@
    Local maxima of HarmonicRatio values occur when acoustic beating is slowest. Therefore, high HarmonicRatio values are associated with low ‘roughness’ and slow acoustic beating.
From this perspective, results involving HarmonicRatio (such as in Section 3.3) may also apply to the period of acoustic beating, i.e., the ‘roughness’ or acoustical ‘grit’.
A.1.6 HarmonicRatio and noise
The tones used to produce the results in Figures 1619 were composed of a finite number of sinusoidal components and did not contain noise. As mentioned earlier with reference to Schneider and Frieler, (2009), inharmonicity can also arise from noise. In this case, low HarmonicRatio values can be understood as resulting from inharmonic relations between the (infinite set of notional) sine waves constituting the noise, whose frequencies can differ infinitesimally.
From the perspective of discrete spectral transforms, an infinite-length signal can be identified as noise-free when spectral bins with non-zero values are separated by bins with zero values. Conversely, noise may be characterised by several contiguous bins with non-zero values. However, defining ‘noise’ in a spectral transform is challenging for at least the following two reasons:
@@ A.1.6 HarmonicRatio and noise
    Even the spectrum of a harmonic tone is not a true line spectrum due to its finite duration.
@@
    A transform occurs over an audio window, and a sine wave with slight frequency changes within the window (e.g., from vibrato) will produce a frequency distribution, not a line.
@@
As it is difficult to distinguish noise from spectral lines in practice, we choose not to distinguish between them. We fall back on the objective information we have. Thus, from a spectral domain perspective, the signal’s information is represented as values attached to spectral bins. Frequency intervals are only defined between the spectral bin center frequencies, and any information we derive from the signal originates from the values attached to the spectral bins. From this point of view, a feature representing the ‘degree of inharmonicity’ in the signal may be conceptualised as deriving from the ensemble of pairwise combinations between all frequency bins (positions and energy values). As a result, a feature representing the ‘degree of inharmonicity’ in the signal will not be able to distinguish between ‘inharmonic sounds which have little if any relevance for music (e.g., white or pink noise)’ (Schneider and Frieler,, 2009) and ‘coherent’ inharmonic signals, which ‘sound as stable and smooth as harmonic signals’ (De Boer,, 1956), i.e., inharmonic complex tones with a finite number of sinusoidal components.
In light of the relation between HarmonicRatio and pitch strength mentioned at the beginning of Section 3.2.1, the ‘first peak of the auto-correlation function’ (Patterson et al.,, 1996; Yost,, 1996, 1997; Shofner and Selas,, 2002) is unable to determine whether a lower pitch strength results from relations between partials, or from noise. We therefore need to find a method to estimate the relative extent to which inharmonicity results from the interaction between (approximately) discrete partials as opposed to noise.
@@
Figure 20: HarmonicRatio, inharmonicity and noisy signals. (a) The signals corresponding to the two spectra have the same HarmonicRatio value of 0.9558. The red signal is inharmonic, with an inharmonicity coefficient of 0.002 according to Fletcher et al., (1962). The blue signal is the sum of a harmonic complex tone and white noise. The white noise’s RMS power is 13.5dB below that of the harmonic tone. (b) Comparative decrease of HarmonicRatio values starting from a sine wave. The blue curve is the result of added white noise. The red curve is the result of adding sine waves of random frequencies. The frequencies being random, the experiment is repeated 200 times. The red area shows the 25th and 75th percentiles of the results.
@@
Figure 20 (a) shows how an inharmonic complex tone and a harmonic complex tone with white noise can produce the same HarmonicRatio value. Conversely, for a HarmonicRatio value less than 1, the original audio may either be a non-inharmonic tone with added noise or a noiseless inharmonic complex tone.
@@
Figure 20 (b) illustrates how a sum of sine waves and a sine wave with white noise can result in the same HarmonicRatio values. The graph shows that a sine wave superposed with noise of equal RMS power is, on average, as inharmonic as a sum of 14 sine waves with the same energy and random frequencies.
A.1.7 Inharmonicity as 1--HarmonicRatio.
@@
The output of the HarmonicRatio feature reflects the pairwise relations between values in all spectral bins, whether these values originate from noise or partials. High HarmonicRatio values indicate harmonic partials without noise. Any deviation lowers the HarmonicRatio values. This deviation can come from inharmonicity in complex tones, inharmonic relations from certain combinations of complex tones, or inharmonic relations arising from contiguous non-zero spectral bins, which may be due to noise. This summary defines an ‘inharmonicity’ feature, termed ‘HR-inharmonicity’, directly derived from HarmonicRatio:
@@
A.2 ‘Noisiness’, spectral flatness and peak prominence
In the previous section, we stated that:
@@
    both a harmonic complex tone with noise and a noiseless inharmonic complex tone may lead to the same value of HarmonicRatio; and that
@@
    both a discrete ensemble of sine waves with no harmonic relations and white noise added to a sine wave may lead to the same value of HarmonicRatio.
@@
In order to discriminate between noiseless and noisy signals leading to the same value of HarmonicRatio, we need a feature that measures the noisiness of the signal. To that end, we introduce a new metric derived from the AudioSpectrumFlatnessType (ISO,, 2001) / ‘spectral flatness’ (Peeters,, 2004).
@@
Peeters, (2004, p. 20) defines spectral flatness as ‘a measure of the noisiness […] / sinusoidality of a spectrum’. It is calculated as the ratio of the geometric mean to the arithmetic mean, equivalent to the Wiener entropy (WE) of the energy spectrum. A comparison between the WE and the variance of power spectra from the BEA dataset shows a high correlation (Pearson correlation = 0.9973). Thus, in this study, spectral flatness can be understood as similar to the variance of the power spectrum. Intuitively, low spectral flatness values (high WE and high variance) indicate spectral peaks, while high values (low WE and low variance, a ‘flat’ spectrum) indicate a noise-like signal.
@@
Spectral flatness is available in the Essentia toolbox (Bogdanov et al.,, 2013) as FlatnessDB [^5] https://essentia.upf.edu/reference/std_FlatnessDB.html and in the MPEG-7 standard as AudioSpectrumFlatnessType (ISO,, 2001, p. 26). The MPEG-7 standard uses a 1/4 octave logarithmic frequency resolution to evaluate the spectrum. We prefer a higher resolution (1/4 of a semitone) to better define spectral peaks. This interval, close to a Pythagorean comma, aligns with the minimum perceivable pitch difference (Zarate et al.,, 2012) or ‘just noticeable difference’ (Stern and Johnson,, 2010). This higher resolution is feasible because the analysis uses the Fourier transform rather than the short-term Fourier transform (MPEG-7).
@@
We calculate spectral flatness for the music datasets, pink noise, and white noise. The output is $0.998$ for white noise. For pink noise ($0.84$), the value falls between the BEA dataset ($0.83$) and the orchestra dataset ($0.86$). The difference between white and pink noise values arises from using a logarithmic frequency scale. For white noise, bands have constant frequency width, whereas for pink noise, they decrease. On a linear scale, this is reversed: spectral flatness is sensitive to the global spectral envelope. The similarity in values for pink noise and music undermines the goal of measuring noisiness. To assess peak prominence (i.e., noisiness) independently of the global envelope, we apply median filtering to the spectrum before calculating WE. The filter uses 8-bin windows (1 tone), with experiments showing comparable results for window widths between 4 and 16 bins.